I need help with this word problem: The scale drawing of a rectangular park uses a scale of 1 centimeter= 74 meters. The drawing is 11 centimeters by 18 centimeters. What are the actual dimensions of the park?
Answer by jim_thompson5910(35256)

1 centimeter = 74 meters

11*1 centimeter = 11*74 meters

11 centimeters = 814 meters


So the actual width is 814 meters

---------------------------------------------------
1 centimeter = 74 meters

18*1 centimeter = 18*74 meters

18 centimeters = 1332 meters


So the actual length is 1332 meters

=============================================================
Answer:

So in short, the actual length of the park is 1332 meters and the actual width of the park is 814 meters.